OpenStudy (anonymous):

How many hotdogs and how many cans of soda were sold? Mitchell works at a concession stand selling hotdogs and cans of soda. A hotdog costs $3.00 and a can of soda costs $1.00. One night, he notices that he sold a total of 350 items and collected a total of $624.

TheSmartOne (thesmartone):

hotdogs=h cans of soda=s

TheSmartOne (thesmartone):

s+h=350 3h+1s=624

TheSmartOne (thesmartone):

s=350-h 3h+1(350-h)=624 3h+350-h=624

TheSmartOne (thesmartone):

2h=974 h=487

OpenStudy (anonymous):

So in order to find out how many cans of soda, would i input the amount of hotdogs? @TheSmartOne

TheSmartOne (thesmartone):

Yes are you sure that the question is correct. If you did that then s=-137

TheSmartOne (thesmartone):

@ganeshie8 @iambatman @e.mccormick

TheSmartOne (thesmartone):

s is not supposed to be negative

OpenStudy (anonymous):

That's why I was so confused.... Here is the exact question.... @TheSmartOne

TheSmartOne (thesmartone):

@Hero @ganeshie8 @iambatman

TheSmartOne (thesmartone):

@Reemxoxo you should bump your question so more people can help you.

OpenStudy (anonymous):

Okay, i'll have to repost it